OpenStudy (anonymous):

I'll give you a metal , help ...... :) One of the factors of the polynomial x³ + 5x^2 + 6x is (x^2+ 3x). What is the other factor? a . x-2 b.x+2 c.x+1 d.x-1

OpenStudy (anonymous):

\[x^3+5 x^2+6 x=x (2+x) (3+x)=(2+x)\left(3 x+x^2\right) \]

OpenStudy (anonymous):

b.

OpenStudy (anonymous):

Thank you for the medal.

OpenStudy (anonymous):

no thank you (: